Karin biked at 13 mph for a time but got a flat tire. She then walked at 5 mph. She traveled a total of 93 miles. Had she biked the whole time, she would have gone 117 miles. How many hours did she walk?

Answer by checkley71(8403)

117/13=9 hours if she biked all the way.
13*x+5(9-x)=93
13x+45-5x=93
8x=93-45
8x=48
x=48/8
x=6 hours she biked.
9-6=3 hours for her walking time.